(8.0*10exp3)(2.0*10exp2) I am not remembering how to solve these equations. I thought you do what is in the () first, then multiply the two answers together. Then I need to right the answer in scientific notation which I am not understanding how to do.
Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

(8.0*10^3)(2.0*10^2)
= (8*2)(10^3 * 10^2)
= 16 x 10^5
= 1.6 x 10^6
================
Cheers,
Stan H.

Answer by scott8148(6628)

it is all multiplication

8.0*10^3*2.0*10^2=16.0*10^5

scientific notation means one non-zero digit to the left of the decimal point
__ 16.0 becomes 1.6 __ since it got smaller, 10^5 becomes 10^6 to maintain the value

1.6*10^6